OpenStudy (flygirlmandi):

-0.6n is less than or equal to 9.6 ? what is n ? a=n is greater than or equal to -16 b=n is greater than or equal to 16 c=n is less than or equal to -16 d=n is less than or equal to 16 thanks :)

imqwerty (imqwerty):

okay so we have this- \(-0.6n \le 9.6\) and now we gotta comment on value of n just divide both sides by \(-0.6\) but remember that the sign of inequality will reverse when you divide by a negative quantity :)

OpenStudy (flygirlmandi):

\[1n \le -16 \] so is this correct ?

OpenStudy (lynfran):

no the sign needs to be reverse

OpenStudy (flygirlmandi):

\[1n \ge -16 \] this is ?

OpenStudy (lynfran):

yea

OpenStudy (lynfran):

the first option

OpenStudy (flygirlmandi):

thanks ! :)

OpenStudy (lynfran):

welcome